OpenStudy (anonymous):

can someone explain how -2y is the cube root of -8y^3

jimthompson5910 (jim_thompson5910):

Take note that 2^3 = 2*2*2 = 8 So (-2)^3 = (-2)*(-2)*(-2) = -8 Also, y^3 = y*y*y = y^3 So (-2y)^3 = (-2y)*(-2y)*(-2y) = -8y^3 The cube root simply "undoes" the cubing.

OpenStudy (anonymous):

-2y times three = 8y^3 (-2y)*(-2y)*(-2y) = (4y^2)*(-2y) = -8y^3